A= π/3 and B π/3, then what is the value of sin (A+B)

 

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
A = \pi/3 and
B = \pi/3
 
Hence
Sin (A +B) = sin (2\pi/3)
 = sin(120°) = cos 30° = (\sqrt3)/2
 
 
Regards
Arun (askIITians forum expert)